OpenStudy (anonymous):

if you are driving at 65 mph and the town you are traveling to is 500 miles away, how many hours will it take?

OpenStudy (anonymous):

Time = Distance Divided by Speed = 500 miles/ 65 mph = 100/13 hours = 7.7 hours Approx

OpenStudy (lstrasmann):

\[D = rt\]\[500 = 65x\]\[\left( 500 \over 65 \right) = \left( 65x \over 65 \right)\]\[7.692307.. = x\]so.. approx. 7 hrs 42 min